# DPmoire: A tool for constructing accurate machine learning force fields in moiré systems

### v1

@@ I Introduction
To address this computational challenge, researchers have developed parameterized continuum models that are better suited for structures with small twist angles  [23, 24, 25, 26, 27, 28, 29]. While these models provide a computationally feasible alternative, they typically do not reach the accuracy levels of DFT relaxation. For materials such as graphene [30, 31, 32] and transition metal dichalcogenides (TMDs) [33], empirical force fields have been effectively utilized for structural relaxation [34, 35, 36, 12, 37]. However, in other systems, robust and extensively validated empirical potentials remain scarce, limiting the scope of studies that can be conducted.
@@ III Results
Then, the MLFF is constructed utilizing the previously determined optimal vdW corrections for both AA and AB stacking configurations of MX2 (M = Mo, W; X = S, Se, Te) materials, as thoroughly discussed in Appendix. B. We specifically examined AA WSe2 and AA MoS2 as representative examples. The efficacy of the MLFF is demonstrated through a comparison of predicted and DFT-calculated forces within the test set, as illustrated in Fig. LABEL:fig:Ferror. The comparison shows a strong alignment between the MLFF predictions and the DFT calculations, with root mean square errors of 0.007 eV/Åand 0.014 eV/Åfor WSe2 and MoS2, respectively, underscoring the accuracy of the MLFF in capturing the essential physical interactions in these materials.
@@ Acknowledgements.
This work was supported by the Science Center of the National Natural Science Foundation of China (Grant No. 12188101), the National Key R&D Program of China (Grant No. 2023YFA1607400, 2024YFA1408400, 2022YFA1403800), the National Natural Science Foundation of China (Grant No.12274436, 11925408, 11921004), and H.W. acknowledge support from the New Cornerstone Science Foundation through the XPLORER PRIZE.

### v2

@@ I Introduction
To address this computational challenge, researchers have developed parameterized continuum models that are better suited for structures with small twist angles  [23, 24, 25, 26, 27, 28, 29, 30]. While these models provide a computationally feasible alternative, they typically do not reach the accuracy levels of DFT relaxation. For materials such as graphene [31, 32, 33] and transition metal dichalcogenides (TMDs) [34], empirical force fields have been effectively utilized for structural relaxation [35, 36, 37, 12, 38]. However, in other systems, robust and extensively validated empirical potentials remain scarce, limiting the scope of studies that can be conducted.
@@ III Results
Then, the MLFF is constructed utilizing the previously determined optimal vdW corrections for both AA and AB stacking configurations of MX2 (M = Mo, W; X = S, Se, Te) materials, as thoroughly discussed in Appendix. B. We specifically examined AA WSe2 and AA MoS2 as representative examples. The efficacy of the MLFF is demonstrated through a comparison of predicted and DFT-calculated forces within the test set, as illustrated in Fig. 4. The comparison shows a strong alignment between the MLFF predictions and the DFT calculations, with root mean square errors of 0.007 eV/Åand 0.014 eV/Åfor WSe2 and MoS2, respectively, underscoring the accuracy of the MLFF in capturing the essential physical interactions in these materials.
@@ Acknowledgements.
J.L. thank Hanqi Pi for helpful discussions. This work was supported by the Science Center of the National Natural Science Foundation of China (Grant No. 12188101), the National Key R&D Program of China (Grant No. 2023YFA1607400, 2024YFA1408400, 2022YFA1403800), the National Natural Science Foundation of China (Grant No.12274436, 11925408, 11921004), and H.W. acknowledge support from the New Cornerstone Science Foundation through the XPLORER PRIZE.